OpenStudy (anonymous):

V= s^2 + 1/2sh, for h

OpenStudy (anonymous):

V=s^2+1/2 sh V-s^2=1/2 sh sh=2V-2s^2 h=2V/s - 2s

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Can you break it down more please

OpenStudy (anonymous):

At first I added -s^2 on both sides

OpenStudy (anonymous):

got that step

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then after that I multiplied both sides by 2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

V-s^2=1/2 sh 2V-2s^2=sh sh=2V-2s^2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Why do you switch the equal sign around?

OpenStudy (anonymous):

To get h in the left hand side

OpenStudy (anonymous):

Oh okay I get it

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Welcome